# INSIGHT: End-to-End Neuro-Symbolic Visual Reinforcement Learning with Language Explanations

### v1

@@ 4.1 Experimental Setup
Tasks  We consider the online RL setting and select nine Atari tasks: Pong, BeamRider, Enduro, Qbert, SpaceInvaders, Seaquest, Breakout, Freeway, and MsPacman. They range from simple motion control (e.g. Pong) to complex decision-making (e.g. Seaquest), and cover issues such as clean background (e.g. SpaceInvaders) vs noisy background (e.g. BeamRider) and constant object (e.g. Pong) vs varying objects (e.g. Enduro). So with them we can examine INSIGHT comprehensively. When analyzing design choices, we report results on five of them due to the space limit.
@@
Baselines  For task performance, INSIGHT is compared with Neural, DSP, Diffses, NUDGE, and CGP. Neural is a deep RL alternative for INSIGHT that uses the same network architecture but does not learn from the frame-symbol datasets. The remaining four are approaches for NS-RL. DSP (Landajuela et al., 2021) searches for symbolic policies from low-dimensional input using a recurrent neural network, while Diffses (Zheng et al., 2022) uses frozen SPACE models to extract states and search for symbolic policies using genetic programming. NUDGE (Delfosse et al., 2023a) relies on the ground-truth structured states and represent policies with first-order differentiable logic. Lastly, CGP (Wilson et al., 2018) leverages cartesian genetic programming for learning policies from images.
@@
Table 2: Inference time per step on Pong. The testing conditions are provided in Sec. A.4. INSIGHT is as efficient as Neural are is an order of magnitude faster than SA-Neural and SPACE-Neural.
@@
Figure 3: Performance of INSIGHT and its variants for five tasks. Refer to Fig. A1 for results for the remaining four tasks. Test returns are normalized so that INSIGHT corresponds to one and random policy is zero. Fixing the perception module during policy learning (i.e., Fixed) hinders performance for all tasks, indicating that it is crucial to refine states with reward signals. The results of w/o Pretrain and w/o NG show that pre-training the perception module and the proposed neural guidance scheme also contribute to the performance of INSIGHT.
4.2 Results
Task Performance  Tab. 1 shows that except for Pong, where visual observations are less involved, NS-RL baselines are outperformed by Neural by a wide margin. In contrast, INSIGHT can match or even outperform Neural. In addition, in Tab. A8 we show that our implementation for Neural matches the performance of CleanRL (Huang et al., 2022), an open-source implementation for deep RL agents. These results highlight that INSIGHT, by learning structured states and symbolic policies simultaneously, overcomes the performance drawback of existing NS-RL approaches. We now present an analysis for learning object coordinates and symbolic policies to explain the improvement.
Table 3: MAE and F-MAE for coordinate prediction. The lower, the better. F-MAE measures the mean absolute error of coordinate prediction for policy-related objects. Numbers are multiplied by 100. Refer to Tab. A7 for results for the other four tasks. By comparing the results of INSIGHT with those of Fixed, we see that refining state representations with rewards improves coordinate prediction for policy-related objects. The proposed neural guidance scheme is crucial for such improvement, as implied by results of w/o NG.
@@ 4.2 Results
Learning Object Coordinates  As mentioned in Sec. 3.1, the perception module of INSIGHT is more efficient than unsupervised tracking models such as SPACE. This argument is well supported by Tab. 2, which shows that both SPACE-Neural and SA-Neural are an order of magnitude slower than INSIGHT. In other words, for 10 million steps, the total inference time of SPACE-Neural exceeds 138 hours, whereas that of INSIGHT is less than six hours. In addition, we show in Tab. A5 that agents learn faster when learning object coordinates from the frame-symbol datasets. A possible reason is that the frame-symbol dataset provides direct supervision for object locations, while unsupervised methods are trained to reconstruct whole images and thus less focused on individual objects.
@@
Figure 4: Influence of the weight for sparsity regularization $\lambda_{\text{reg}}$, the width/layer of the EQL actor, and the weight of $\mathcal{L}_{\text{cnn}}$ for SpaceInvaders. Results for other tasks are provided in Fig. A2. The performance of INSIGHT is quite robust against hyper-parameters.
@@
Our main claim in this paper is that it is essential to refine structured states with both visual and reward information. We now examine this claim from the perspective of task performance and coordinate prediction. Fig. 3 shows that when compared to the variant Fixed, INSIGHT has significant better performance for all five task, demonstrating that INSIGHT’s performance is largely determined by its ability to refine structured states with reward signals. Tab. 3 further reveals a clue for the performance improvement. Compared to Fixed, INSIGHT has higher MAE for four tasks but lower F-MAE for all five tasks. In addition, readers may refer to Fig. A3 for visualizations of predicted coordinates. These findings suggest that reward signals can guide agents to improve the coordinate prediction of policy-relevant objects.
@@
In Sec. 3.1, we suggest pre-training the perception module before policy learning. Fig. 3 shows that the pre-training step indeed improves performance for four tasks, and Tab. 3 indicates that the improvement may be the result of better prediction for policy-relevant objects. Meanwhile, this extra step does complicates the training protocol of INSIGHT. One possible improvement is to use off-policy algorithms for policy learning, which allow the perception module to be optimized for more gradient steps.
@@
Learning Symbolic Policies  We introduce the neural guidance scheme to address the limited expressiveness of object coordinates. The effectiveness of this scheme is supported by results in Fig. 3–w/o NG performs worse than INSIGHT for three tasks out of five tasks. Moreover, Coor-Neural is outperformed by Neural for six tasks in Tab. 1, suggesting that the object coordinates are not sufficiently expressive for online policy learning. Tab. 3 reveals that unlike the case of INSIGHT, the F-MAE and MAE of w/o NG are almost the same, implying that w/o NG cannot improve coordinate prediction using reward signals. Thus, the neural guidance scheme plays a key role in improving performance by refining coordinate prediction with rewards.
@@
The Influence of Hyper-parameters  Finally, Fig. 4 shows the influence of four hyper-parameters for SpaceInvaders. Results for other tasks are provided in Fig. A2. It turns out that the performance of INSIGHT are quite robust against the hyper-parameters. Since our goal is to illustrate the efficacy of INSIGHT rather than benchmarking, we use the same values for hyper-parameters for all tasks.
@@ 4.3 Language Explanations
The left part of Fig. 5 presents interpretations of a policy learned for Pong. Readers may refer to the bottom right of Fig. 5 for screenshots of this game. Like a ping-pong game, an agent controls the right paddle to hit the ball to its opponent’s side and scores a point if the opponent fails to hit the ball back. Both paddles move only vertically.
@@
The right part of Fig. 5 showcases decision explanations. The LLM correctly deduces the motion of the ball and the opponent’s paddle from the input coordinates, but it makes a mistake for the agent’s paddle, possibly due to fluctuations in the coordinate predictions. Moreover, by inspecting the gradients, it is pointed out that the decision is more sensitive to the position of the opponent’s paddle than the ball, suggesting that the agent is exploiting the opponent’s fixed policy. In the conclusion, the interpretations (e.g., to intercept or hit the ball) are supported by facts (e.g., the opponent’s motion), which can improve their credibility.
Overall, both the policy interpretation and the decision explanation are constructed from the information provided and are free of factual errors. They reveal some patterns in the agent’s decision-making process, although there are minor mistakes. Nevertheless, they suffice to explain the learned policies to users who are not familiar with NS-RL.
@@ 5 Conclusions
We propose INSIGHT, a framework that uses object coordinates as structured state representations and learns symbolic policies from visual input. INSIGHT is able to refine the structured states with reward signals by distilling vision foundation models into a scalable perception module, and it leverages a new neural guidance scheme to learn competitive symbolic policies from object coordinates, thereby overcoming the performance bottleneck of previous NS-RL approaches. Moreover, to improve model transparency for non-expert users, INSIGHT can generate language explanations for learned policies and specific decisions with LLMs. With experiments on nine Atari tasks, we show that INSIGHT outperforms all NS-RL baselines, and the improvement can be explained by the improved coordinate prediction for policy-relevant objects. We also showcase language explanations for learned policies and decisions.
@@
While the nine Atari tasks have already covered several critical issues, additional experiments for continuous control or realistic observation can complement our evaluation. Another limitation of INSIGHT is that EQL networks cannot express logical operations required by some reasoning tasks. Finally, quantitative evaluation of the explanations is also an interesting future topics.
@@ A.1 Details of the Frame-Symbol Dataset
The generation of the dataset begins with a comprehensive training regimen of 10 million steps using the neural baseline. This regimen incorporates a neural policy tailored for interaction with the environment. During the final 1 million steps, specifically starting from the 9-millionth step, we capture a total of 10,000 frames, thereby forming an unsupervised frame dataset. This phase primarily focuses on documenting environmental interactions via image captures, with a special emphasis on acquiring objects in high-reward scenarios.
@@
Following this, the initial frame undergoes processing through FastSAM, which identifies and segregates 256 unique objects by extracting their masks. These masks serve as inputs for the DeAot module, enabling object tracking across the subsequent 9,999 frames. FastSAM re-evaluates the scene every tenth frame to include new objects, leading to an increase in the object count. This increment prompts DeAot to start tracking these newly identified objects. For improved segmentation and tracking, all frames are resized to a resolution of 1024×1024 pixels.
@@
During the segmentation phase using FastSAM, objects are excluded if their confidence level does not meet or exceed a threshold of 0.9. This stringent selection criterion is pivotal for minimizing misidentifications attributable to environmental factors. Moreover, within the same frame, a distinction is made between connected and non-connected masks. Connected masks that do not overlap with previously tracked objects are deemed new. Conversely, objects that overlap with an existing mask by 50% or less are also classified as new. The rationale for the former is to avoid fragmenting a single object into multiple smaller segments as much as possible, and for the latter, it is because the tracking module tends to recognize non-connected, similar objects as a single entity, necessitating their re-identification and enumeration.
@@
In the dataset’s final development phase, the FastSAM module is deactivated, allowing DeAot to autonomously track objects from the initial to the final frame. The resulting dataset encompasses detailed information like object coordinates, bounding boxes, and RGB values, linked to their corresponding frames. Ultimately, this leads to the formation of a comprehensive frame-symbol dataset, where frames are stored at a resolution of 512×512 pixels. This process mitigates a previously identified limitation where the tracking model struggled to consistently detect objects in each frame. Conforming to the standard supervised training approach, the dataset is divided into training and test sets in an 80:20 ratio.
@@ B.1 Architecture and Hyperparameters
Table A1: Hyperparameters of the CNN Encoder
@@
Table A2: Hyperparameters of EQL Network
@@
Table A3: Hyperparameter of Pretraining
@@
Table A4: Hyperparameters of Policy Learning
@@
In the policy learning phase, following the common settings of Atari tasks, INSIGHT interacts with the environment for 10M steps. PPO with a learning rate of 2.5e-4 uses the collected rewards to optimize the model. The batchsize size of each update is 1024. In order to ensure that the network can learn useful policies while ensuring the coefficient of the policy, we choose to use 1e-3 regularization and increase the regularization coefficient from 0 to 1 in each update. In addition, we multiply the coefficient of Eq. 3 by 2 to enhance the accuracy of coordinate prediction. All detailed parameters are summarized in Tab. A4.
@@ C.1 Comprehensive Evaluation of Representations
Table A5: Performance of SA-Neural, SPACE-Neural, Coor-Neural, and Neural when trained for one million steps. Coor-Neural outperforms SA-Neural and SPACE-Neural on most tasks and is the only method that can compete with Neural, which shows that agents learn faster when using our proposed method of learning structured state representations.
@@
BeamRider $684.8\pm 587.1$ $579.6\pm 348.4$ $\textbf{846.1}\pm 92.2$ $551.6\pm 25.1$
Enduro $0\pm 0$ $8.1\pm 8.4$ $\textbf{45.4}\pm 63.5$ $9.7\pm 6.9$
@@
MsPacman $643.2\pm 58.2$ $60.4\pm 134.4$ $1180.2\pm 245.9$ $\textbf{1289.2}\pm 106.7$
Table A6: Performance of SA-Neural, SPACE-Neural, Coor-Neural, and Neural when trained for three million steps. Coor-Neural consistently achieves higher performance than SA-Neural and SPACE-Neural in the case of convergence.
@@
These methods were trained for only one million steps due to the poor efficiency of SPACE-Neural. Similar experiment setup has been utilized by Delfosse et al. (2023b). We also report results for three million steps in Tab. A6, by which point the performances on Pong and Freeway had stabilized. Note that compared to slot attention or the SPACE model, our perception module leads to better performance. This is because our image-symbol dataset provides direct supervision for the location of objects. On the contrary, unsupervised methods are trained to reconstruct whole images and thus less focused on individual objects.

### v2

@@ 4.1 Experimental Setup
Tasks  We consider the online RL setting and nine Atari tasks: Pong, BeamRider, Enduro, Qbert, SpaceInvaders, Seaquest, Breakout, Freeway, and MsPacman, which widely used in prior works (Delfosse et al., 2023b; Landajuela et al., 2021; Zheng et al., 2022). They range from simple motion control (e.g. Pong) to complex decision-making (e.g. Seaquest), and cover issues such as clean background (e.g. SpaceInvaders) vs noisy background (e.g. BeamRider) and constant object (e.g. Pong) vs varying objects (e.g. Enduro). So with them we can examine INSIGHT comprehensively. In the meantime, results on more realistic tasks are also important for validating NS-RL algorithms, so we also include results on MetaDrive (Li et al., 2022), a challenging environment for autonomous driving.
@@
Table 2: INSIGHT outperforms neural baselines on MetaDrive. This table shows the performance of INSIGHT, Neural and Random on MetaDrive when trained fro 1M, 2M, and 5M steps.
Success Rate INSIGHT Neural Coor-Neural Random
1M $\textbf{0.37}\pm 0.21$ $0.17\pm 0.12$ $0.07\pm 0.05$ $0\pm 0$
2M $\textbf{0.42}\pm 0.13$ $0.19\pm 0.06$ $0.21\pm 0.14$ $0\pm 0$
5M $\textbf{0.63}\pm 0.11$ $0.49\pm 0.11$ $0.41\pm 0.12$ $0\pm 0$
Baselines  For task performance, INSIGHT is compared with Neural, DSP, Diffses, NUDGE, and CGP. Neural is a deep RL alternative for INSIGHT that uses the same network architecture but does not learn from the frame-symbol datasets. The remaining four are approaches for NS-RL. DSP (Landajuela et al., 2021) searches for symbolic policies from low-dimensional input using a recurrent neural network, while Diffses (Zheng et al., 2022) uses frozen SPACE models to extract states and search for symbolic policies using genetic programming. NUDGE (Delfosse et al., 2023b) relies on the ground-truth structured states and represent policies with first-order differentiable logic. Lastly, CGP (Wilson et al., 2018) leverages cartesian genetic programming for learning policies from images.
@@
Table 3: INSIGHT is as efficient as Neural. This table shows the inference time per state on Pong. The testing conditions are provided in Sec. A.4. INSIGHT is as efficient as Neural are is an order of magnitude faster than SA-Neural and SPACE-Neural.
@@
Figure 3: Each component of INSIGHT is critical for overall performance. Detailed performance analyses of INSIGHT and its variants across five tasks are presented. Refer to Fig. A1 for results for the remaining four tasks. Test returns are normalized so that INSIGHT corresponds to one and random policy is zero. Fixing the perception module during policy learning (i.e., Fixed) hinders performance for all tasks, indicating that it is crucial to refine states with reward signals. The results of w/o Pretrain and w/o NG show that pre-training the perception module and the proposed neural guidance scheme also improves performance.
Table 4: INSIGHT significantly enhances coordinate prediction for policy-relevant objects through policy learning. MAE and F-MAE measure the coordinate prediction error, with lower values indicating better performance. F-MAE measures the mean absolute error of coordinate prediction for policy-related objects. Numbers are multiplied by 100. Refer to Tab. A7 for results for the other four tasks. By comparing the results of INSIGHT with those of Fixed, we see that refining state representations with rewards improves coordinate prediction for policy-related objects. The proposed neural guidance scheme is crucial for such improvement, as implied by results of w/o NG.
@@
4.2 Results
Improved Task Performance  Tab. 1 shows that except for Pong, where visual observations are less involved, NS-RL baselines are outperformed by Neural by a wide margin. In contrast, INSIGHT can match or even outperform Neural. In addition, in Tab. A8 we show that our implementation for Neural matches the performance of CleanRL (Huang et al., 2022), an open-source implementation for deep RL agents. Tab. 2 shows the success rates of INSIGHT and neural baselines over 1M, 2M, and 5M timesteps for Metadrive. INSIGHT outperforms the neural baselines at all timesteps. These results highlight that INSIGHT, by learning structured states and symbolic policies jointly, overcomes the performance drawback of existing NS-RL approaches. We now present an ablation analysis to explain the improvement.
@@ 4.2 Results
Better Efficiency  As mentioned in Sec. 3.1, the perception module of INSIGHT is designed to be efficient. Tab. 3 shows that INSIGHT is an order of magnitude faster than both SPACE-Neural and SA-Neural, which verifies this design. In addition, Tab. A5 shows that agents learn faster when learning from $\mathcal{D}_{\text{symbol}}$. A possible reason is that $\mathcal{D}_{\text{symbol}}$ provides direct supervision for object locations. When trained with image reconstruction objectives, agents are not forced to focused on individual objects.
@@
Figure 4: INSIGHT demonstrates robustness to hyper-parameters. Examining the influence of sparsity regularization weight $\lambda_{\text{reg}}$, the EQL actor’s width/layers, and the weight of $\mathcal{L}_{\text{cnn}}$ on SpaceInvaders performance. Results for additional tasks are available in Fig. A2. Overall, INSIGHT shows substantial robustness to variations in hyper-parameters.
@@
Benefits of End-to-End Policy Learning  We now examine our claim for refining structured states with both visual and rewards from the perspective of task performance and coordinate prediction. Fig. 3 shows that when compared to the variant Fixed, INSIGHT has significant better performance for all five task, demonstrating that INSIGHT’s performance is largely determined by its ability to refine structured states with reward signals. Tab. 4 further reveals a clue for the performance improvement. Compared to Fixed, INSIGHT has higher MAE for four tasks but lower F-MAE for all five tasks. In addition, readers may refer to Fig. A3 for visualizations of predicted coordinates. These findings suggest that reward signals can guide agents to improve the coordinate prediction of policy-relevant objects.
@@
Benefits of Pre-Training the Perception Module  In Sec. 3.1, we suggest pre-training the perception module before policy learning. Fig. 3 shows that the pre-training step indeed improves performance for four tasks, and Tab. 4 indicates that the improvement may be the result of better prediction for policy-relevant objects. Despite the imperfections and redundancy in the representation space obtained through pre-training, it still offers valuable information about policy-related objects, serving as a strong initialization for subsequent end-to-end training. Meanwhile, this extra step does complicate the training protocol of INSIGHT. One possible improvement is to use off-policy algorithms for policy learning, which allow the perception module to be optimized for more gradient steps.
@@
Efficacy of the Proposed Neural Guidance Scheme  We introduce the neural guidance scheme to address the limited expressiveness of object coordinates. The effectiveness of this scheme is supported by results in Fig. 3–w/o NG performs worse than INSIGHT for three tasks out of five tasks. Moreover, Coor-Neural is outperformed by Neural for six tasks in Tab. 1, suggesting that the object coordinates are not sufficiently expressive for online policy learning. Tab. 4 reveals that unlike the case of INSIGHT, the F-MAE and MAE of w/o NG are almost the same, implying that w/o NG cannot improve coordinate prediction using reward signals. Thus, the neural guidance scheme plays a key role in refining coordinate prediction with rewards.
@@
Robustness Against Hyper-Parameters  Lastly, Fig. 4 shows the influence of four hyper-parameters for SpaceInvaders, which indicates that the performance of INSIGHT are quite robust against the hyper-parameters. Results for other tasks are provided in Fig. A2. Since our goal is to illustrate the efficacy of INSIGHT rather than benchmarking, we use the same values for hyper-parameters for all tasks.
@@ 4.3 Language Explanations
The left part of Fig. 5 presents interpretations of a policy learned for Pong, which requires an agent to control the right paddle to hit the ball to the left-hand side. The agent earns a point if the opponent fails to hit the ball back. Both paddles move only vertically.
@@
The right part of Fig. 5 showcases decision explanations. While the movements of the ball is correctly deduced, there is a mistake for the agent’s paddle, possibly due to fluctuations in the coordinate predictions. The generated explanations reveals that the decision is more sensitive to the position of the opponent’s paddle than the ball, suggesting that the agent is exploiting the opponent’s fixed policy. Similar phenomena are also observed in recent NS-RL works (Delfosse et al., 2024), which highlights the importance of inspecting learned policies manually. Nevertheless, INSIGHT can generate credible explanations by associating interpretations with facts.
Overall, despite the minor mistakes, both the policy interpretation and the decision explanation are friendly to non-expert users and reveal some patterns in the agent’s decision-making process.
@@ 5 Conclusions
We propose INSIGHT, a framework that uses object coordinates as structured state representations and learns symbolic policies from visual input. INSIGHT is able to refine the structured states with reward signals by distilling vision foundation models into a scalable perception module, and it leverages a new neural guidance scheme to learn competitive symbolic policies from object coordinates, thereby overcoming the performance bottleneck of previous NS-RL approaches. Moreover, to improve model transparency for non-expert users, INSIGHT can generate language explanations for learned policies and specific decisions with LLMs. With experiments on nine Atari tasks and a MetaDrive task, we show that INSIGHT outperforms all NS-RL baselines and further reveal that the improvement can be explained by the improved coordinate prediction for policy-relevant objects. We also showcase language explanations for learned policies and decisions.
@@
A limitation of INSIGHT is that EQL networks cannot express logical operations required by some reasoning tasks. Furthermore, quantitative evaluation of the explanations is also an interesting future topic.
Acknowledgments
C. Fang was supported by National Science and Technology Major Project (2022ZD0114902), the NSF China (No. 62376008). The authors would like to thank four anonymous reviews for constructive feedback.
@@ A.1 Details of the Frame-Symbol Dataset
The generation of the dataset begins with a comprehensive training regimen of 10 million steps using the neural baseline. This regimen incorporates a neural policy tailored for interaction with the environment. During the final 1 million steps, specifically starting from the 9-millionth step, we captured 10,000 consecutive frames spanning multiple episodes, thereby forming an unsupervised frame dataset. This phase primarily focuses on documenting environmental interactions via image captures, with a special emphasis on acquiring objects in high-reward scenarios.
@@
Following this, the initial frame undergoes processing through FastSAM, which identifies and segregates unique objects by extracting their masks. Without assuming prior knowledge of the number of objects in the environment. To accommodate various environments, we establish a maximum object count of 256 for all experiments, a threshold that is sufficient for all Atari games. The objects are ordered based on their confidence scores from FastSAM. These masks serve as inputs for the DeAot module, enabling object tracking across the subsequent 9,999 frames. To maintain the consistency of object IDs, these 10,000 frames were spliced into a whole video and read by DeAot. FastSAM re-evaluates the scene every tenth frame to include new objects, leading to an increase in the object count. This increment prompts DeAot to start tracking these newly identified objects. For improved segmentation and tracking, all frames are resized to a resolution of 1024×1024 pixels.
@@
During the segmentation phase using FastSAM, objects are excluded if their confidence level does not meet or exceed a threshold of 0.9. This stringent selection criterion is pivotal for minimizing misidentifications attributable to environmental factors. Moreover, within the same frame, a distinction is made between connected and non-connected masks. Connected masks that do not overlap with previously tracked objects are deemed new. Conversely, objects that overlap with an existing mask by 50% or less are also classified as new, i.e., IOU=0.5. The rationale for the former is to avoid fragmenting a single object into multiple smaller segments as much as possible, and for the latter, it is because the tracking module tends to recognize non-connected, similar objects as a single entity, necessitating their re-identification and enumeration.
@@
In the dataset’s final development phase, the dataset was re-segmented starting at $0^{th}$ frame and the FastSAM module is deactivated, allowing DeAot to autonomously track objects from the initial to the final frame. The resulting dataset encompasses detailed information like object coordinates, bounding boxes, and RGB values, linked to their corresponding frames. When an object is missing from a frame, it is noted as ’non-existent’ in the dataset, and its coordinates are set to (0,0). This approach allows for a systematic way to account for absent objects without disrupting the overall tracking and identification process. Ultimately, this leads to the formation of a comprehensive frame-symbol dataset, where frames are stored at a resolution of 512×512 pixels. This process mitigates a previously identified limitation where the tracking model struggled to consistently detect objects in each frame. Conforming to the standard supervised training approach, the dataset is divided into training and test sets in an 80:20 ratio.
@@ B.1 Architecture and Hyperparameters
Table A1: Hyperparameters of the CNN encoder.
@@
Table A2: Hyperparameters of EQL network.
@@
Table A3: Hyperparameter of pretraining.
@@
Table A4: Hyperparameters of Policy Learning.
@@
In the policy learning phase, following the common settings of Atari tasks, INSIGHT interacts with the environment for 10M steps. PPO with a learning rate of 2.5e-4 uses the collected rewards to optimize the model. The batchsize of each update is 1024. In order to ensure that the network can learn useful policies while ensuring the coefficient of the policy, we choose to use 1e-3 regularization and increase the regularization coefficient from 0 to 1 in each update. In addition, we multiply the coefficient of Eq. 3 by 2 to enhance the accuracy of coordinate prediction. All detailed parameters are summarized in Tab. A4.
@@ C.1 Comprehensive Evaluation of Representations
Table A5: Distilled structured representations significantly enhance performance. Evaluating the performance of SA-Neural, SPACE-Neural, Coor-Neural, and Neural after training for one million steps. Coor-Neural surpasses both SA-Neural and SPACE-Neural in most tasks and stands as the only competitor to Neural. This demonstrates the accelerated learning achieved by agents using our proposed method for structured state representation.
@@
BeamRider $684.8\pm 587.1$ $579.6\pm 348.4$ $846.1\pm 92.2$ $551.6\pm 25.1$
Enduro $0\pm 0$ $8.1\pm 8.4$ $45.4\pm 63.5$ $9.7\pm 6.9$
@@
MsPacman $643.2\pm 58.2$ $60.4\pm 134.4$ $1180.2\pm 245.9$ $1289.2\pm 106.7$
Table A6: Distilled structured representations significantly enhance the performance upon convergence. After training for three million steps, the performance of SA-Neural, SPACE-Neural, Coor-Neural, and Neural is evaluated. Coor-Neural consistently outperforms SA-Neural and SPACE-Neural upon convergence.
@@
Distilled Structured Representations Significantly Enhance Performance These methods were trained for only one million steps due to the poor efficiency of SPACE-Neural. Similar experiment setup has been utilized by Delfosse et al. (2023c). We also report results for three million steps in Tab. A6, by which point the performances on Pong and Freeway had stabilized. Note that compared to slot attention or the SPACE model, our perception module leads to better performance. This is because our image-symbol dataset provides direct supervision for the location of objects. On the contrary, unsupervised methods are trained to reconstruct whole images and thus less focused on individual objects.